the perimeter of a rectangle is 140 cm and the ratio of its height to width is 3:4 find the area of the rectangle

Answer by checkley79(3341)

P=2*3M+2*4M
140=6M+8M
140=14M
M=140/14
M=10 THIS IS THE RATIO MULTIPLIER.
3*10=30 FOR THE LENGTH
4*10=40 FOR THE WIDTH.
PROOF:
140=2*30+2*40
140=60+80
140=140